OpenStudy (anonymous):

A square pyramid has a slant height of 13 meters and a height of 12 meters. What is the surface area if the square base has sides of length 10 meters? 360 m2 320 m2 260 m2 120 m2

OpenStudy (anonymous):

my answer is 360

OpenStudy (hba):

^^ Thats the correct answer

OpenStudy (hba):

??

OpenStudy (anonymous):

thanks

OpenStudy (hba):

i am asking you

OpenStudy (anonymous):

yes

OpenStudy (hba):

The surface area of a regular pyramid is equal to the area of its base plus 1/2 times the perimeter of its base times the slant height. SA = 10*10 + (1/2)*(40)(13) SA = 100 + 260 SA = 360 square meters

OpenStudy (anonymous):

then I'm fine